# Learned Kernels for Sparse, Interpretable, and Efficient Medical Time Series Processing

### v3

@@ Abstract
Background:
Rapid, reliable, and accurate interpretation of medical signals is crucial for high-stakes clinical decision-making. The advent of deep learning allowed for an explosion of new models that offered unprecedented performance in medical time series processing but at a cost: deep learning models are often compute-intensive and lack interpretability.
Methods:
We propose Sparse Mixture of Learned Kernels (SMoLK), an interpretable architecture for medical time series processing. The method learns a set of lightweight flexible kernels to construct a single-layer neural network, providing not only interpretability, but also efficiency and robustness. We introduce novel parameter reduction techniques to further reduce the size of our network. We demonstrate the power of our architecture on two important tasks: photoplethysmography (PPG) artifact detection and atrial fibrillation detection from single-lead electrocardiograms (ECGs). Our approach has performance similar to the state-of-the-art deep neural networks with several orders of magnitude fewer parameters, allowing for deep neural network level performance with extremely low-power wearable devices.
Results:
Our interpretable method achieves greater than 99% of the performance of the state-of-the-art methods on the PPG artifact detection task, and even outperforms the state-of-the-art on a challenging out-of-distribution test set, while using dramatically fewer parameters (2% of the parameters of Segade [1], and about half of the parameters of Tiny-PPG [2]). On single lead atrial fibrillation detection, our method matches the performance of a 1D-residual convolutional network, at less than 1% the parameter count, while exhibiting considerably better performance in the low-data regime, even when compared to a parameter-matched control deep network.
Conclusions:
SMoLK models are competitive with deep neural networks for medical time series processing with dramatically fewer parameters. Our method is particularly suited for real-time applications and low-power devices, and it maintains interpretability.
@@ 1 Introduction
Medical signals such as electroencephalograms (EEGs), electrocardiograms (ECGs), and photoplethysmography signals (PPGs) have played a pivotal role in the field of medicine, serving as a crucial medium for diagnosing and monitoring patients’ conditions, facilitating timely interventions, and improving overall healthcare outcomes. These types of signals are usually manually reviewed by trained medical professionals, however, in today’s realm of endless medical data and the popularization of continuous passive monitoring, manual reviewing efforts have become inefficient or even a bottleneck in the patient care process. Traditional techniques to automate signal processing and categorization, while improving speed, often come with limitations in terms of performance. The rise of deep learning yielded a revolution in signal processing, offering models with unparalleled performance (e.g., [3, 4, 5, 6, 7]). However, these models often come with their own set of challenges, particularly in terms of computational demands and a lack of interpretability in their decision-making processes. Indeed, post-hoc saliency methods often fail to correctly identify feature importance in time-series signals [8, 9].
Recognizing the need for high-performing interpretable models, we sought to design a novel architecture that combines the best of both worlds: the performance of deep learning with the computational efficiency and interpretability of traditional methods. We propose a novel method that has performance similar to – and often better than – the state-of-the-art DNN approach with several orders of magnitude fewer parameters. The method learns a set of lightweight flexible kernels to construct a single-layer neural network, providing a new efficient, robust, and interpretable medical signal processing architecture. To train these kernel functions, we introduce parameter reduction techniques: weight absorption absorbs the kernel weighting factors into the kernels themselves and correlated kernel pruning reduces redundancies among the learned kernels.
@@
Our approach achieves compelling results in both the task of detecting artifacts in PPG signals, as well as the detection of atrial fibrillation from single-lead ECGs (an increasingly common task in modern-day wearables). With respect to PPG artifact segmentation, we obtain the best of all worlds: (1) state-of-the-art (SotA) performance and (2) interpretability in a model with (3) several orders of magnitude fewer parameters than the current SotA. For the single-lead ECG task, SMoLK achieves metrics matching a deep ResNet with less than 1% the parameter count, while significantly outperforming deep networks in the low-data regime, even when controlling for overfitting via parameter count. For interpretability (2), our approach yields a tiny-yet-optimized set of convolution kernels that, when applied to the PPG signal and summed following a ceiling function, directly produces a measure of PPG signal quality. This measure can then be thresholded to segment and identify artifacts. When applying these kernels to the signals, the process is “nearly linear” in the sense that the only non-linearity is a ceiling function setting negative values to zero. We can directly and precisely measure the contribution of any individual convolution to the predicted signal quality, and kernel contributions are directly proportional to the output signal. For classification tasks, we can directly invert the process to compute the precise contribution of each part of the signal to the final classification. Our approach is reminiscent of feature-detection-based approaches in that we learn a set of kernels that mimic handcrafted features. For our smallest model, we learn so few kernels (12 total) that we can inspect these features by eye and observe the waveforms learned by our approach with a quick glance. This contrasts with deep neural networks, where there are so many parameters one cannot feasibly inspect the inner workings of the model, and even if one took the time to meticulously inspect these features, it would still be unclear how each feature individually contributes to the output due to the built-in non-linearity.
In terms of efficiency and memory (3), our model reflects that for real-world applications, signal processing can be limited by computational resources or power consumption, as most wearable devices are small computing devices with limited battery life and low compute budget. Furthermore, memory is often a constraint as many wearable devices run multiple processes and monitor several metrics at once. Our models are simple to implement, and easily lend themselves to memory and compute optimization. Our architecture can be quantized easily to 16-bit floating point operations without any appreciable decrease in performance. Our method is lightweight enough to allow real-time assessment of signal quality or classification of arrhythmias with a small enough impact that it could feasibly be implemented as a background process, using only kilobytes of memory.
@@ 2.2 Deep Learning Approaches
Most deep learning methods demand significant computational resources, rendering them impractical for use in wearable devices. An exception is Tiny-PPG [2], which emphasizes model pruning and the development of compact, portable models. Our experiments indicate that we achieve similar performance to Tiny-PPG but with much sparser models (about half the parameters).
Interpreting black box deep learning models is extremely challenging. Instead, most work turns to post-hoc explainability methods, such as the aforementioned GradCAM and SHAP methods [42, 43]. Indeed, even modern explainable approaches to medical signal processing with transformer architectures [45] fail to yield saliency maps that properly attribute features to their classification. In the aforementioned work, saliency maps for ST-elevation myocardial infarction correctly assign high saliency to some elevated ST segments, yet ignore equally prevalent ST segments seemingly at random. Furthermore, an area of equally intense saliency is assigned to a completely blank area of the ECG strip.
3 Datasets and Preprocessing
We train our model on the PPG-DaLiA [46] training split, and evaluate our approach on the same datasets used by the current state-of-the-art algorithms [1, 2], namely PPG-DaLiA [46] test split, WESAD [47], and TROIKA [48]. The PPG-DaLiA dataset is composed of recordings from 15 subjects performing real-world tasks. The data include electrocardiogram (ECG), accelerometer, and electrodermal recordings for various settings/activities, such as sitting still, walking up/down stairs, playing sports (soccer, cycling, walking), and driving. The WESAD data was recorded from wrist and chest PPGs from 15 subjects aged 21-55 (median 28) primarily focusing on varying the emotional state of the subject (neutral, stressed, amused). ECG, accelerometer, electrodermal, and wrist PPG signals were recorded. Lastly, the TROIKA data was recorded from subjects running on a treadmill (ages 18-35). Accelerometer, ECG, and PPG signals were recorded. This dataset represents the most challenging dataset for generalization among the datasets we test in this work, since the subjects were under intense physical demands with poor signal quality, owing to the large amounts of movement.
4 Methods
@@ 4.1 Overview
Figure 1: The processing pipeline for the PPG signal quality segmentation task during training and during inference. First, a PPG signal is normalized to a unit normal by subtracting the mean and dividing by the standard deviation. Next, a set of convolutions is applied. Our smallest model is lightweight enough that all of the convolved signals are displayed in this figure. The convolved signals are ceiled, weighted, and summed, and a sigmoid is applied. Finally, the output is smoothed and thresholded to segment the signal. After training, similar kernels are ablated and kernel weights are absorbed to reduce parameter count.
To achieve this, we learn a set of $M$ convolution kernels of varying numbers to evaluate parameter scaling laws: $M=12$ (small), $72$ (medium), $384$ (large). The $M$ kernels are divided into sets of $M/3$ kernels of short (1.0 seconds), moderate (1.5 seconds), and long (3.0 seconds) sizes, with the goal of learning features of different lengths to improve classification quality. The kernels are convolved with the input PPG signal and the values are ceiled via $\max(0,x)$ to obtain a non-negative signal quality indicator. Additionally, a bias is added to each convolution channel. We compute a weighted sum of the convolutions to produce a single channel that represents a measure of signal quality, which is then transformed to range [0.0, 1.0] via the sigmoid function. We apply a 3rd-order Savitzky-Golay filter ($\sim$0.8-second filter window) to smooth the output, then apply thresholding at a value of 0.5 to produce a binary output segmentation. This pipeline is shown in Figure 1. The addition of a post-processing smoothing filter (Savitzky-Golay in our case) yields smoother segmentation maps and relies on the assumption that PPG artifacts tend to occur in groups, rather than as fragmented artifacts. However, this post-processing filter is not strictly necessary to maintain good DICE metrics and is an optional hyperparameter to be chosen for the use case (see Appendix B).
@@
Since our model consists of a single layer of convolutions, we can directly inspect kernels and measure their contribution to the signal, allowing us to simply prune kernels that detect similar features thereby reducing parameter counts. Additionally, because our model is linear aside from the $\max$ and sigmoid function, we can reduce parameter counts at inference time by merging parameters in a linear fashion. We will describe these novel techniques in more detail below.
4.2 ECG Arrhythmia Classification
@@ 4.2.1 Overview
$\mathbf{f_{m}}$ $=\text{max}(0,\mathbf{x}*\mathbf{k}_{m}+b_{m}).$
@@
$\phi_{m}$ $=\text{Mean}(\mathbf{f_{m}})$
@@
$z_{j}$ $=\sum_{m=0}^{M}w_{mj}\phi_{m}+\sum_{\hat{f}\in\mathcal{F}}w_{\hat{f}j}\hat{f}+b_{j}$
Where $\hat{f}$ are the frequency bands in the power spectrum $\mathcal{F}$, and $w_{\hat{f}j}$ is the corresponding weight of that frequency band for class $j$. Finally, a non-linear activation function can be applied to this logit to generate a probability distribution. We use SoftMax in our case. An illustrated overview of this pipeline can be seen in Figure 2a. We found that a 384-kernel model performs competitively with a 16-layer ResNet (architectural and training details can be found in Appendix A).
Figure 2: a. The pipeline for processing a single lead ECG. First, a set of learned convolutions is applied to the ECG to produce several feature maps. Then, feature-wise means are computed, which are used as inputs to a linear model. Global frequency information, obtained from a power spectrum, are also used as inputs for the linear model to yield a logit. b. The reverse pipeline for interpreting a prediction from the learned kernel model. First, feature maps are generated by the learned kernel convolutions, in the same manner as a. However, instead of computing the mean value of this feature map, the feature maps are instead multiplied element-wise with their corresponding class weights. Lastly, a transposed convolution is applied to compute the importance of each part of the input signal. Notably, this method correctly labels the increased PR-interval as the most important feature in the classification of this ECG as $1^{st}$-degree heart block.
@@ 4.2.2 Interpretability
Because our model is, in essence, a linear combination of feature map mean activations, and each activation was generated from a convolution over the signal, we can directly compute the contribution of each part of the input signal to a given output class by reversing the convolution. We thus compute the contribution map as follows, in Algorithm 1. At a high level, a forward pass is performed as usual and the feature map values are computed. Then, the feature maps are multiplied element-wise with the feature weight for the class in question. Lastly, this product is convolved backward into a contribution map buffer of equal dimension to the input ECG. This allows us to compute exactly how much any given portion of the input signal contributed to the output classification. An illustration of this process can be seen in Figure 2b.
@@ 4.3 Weight Absorption
The quantity $\mathbf{k}_{m}\cdot|w_{m}|$ becomes the new kernel, and the quantity $b_{m}\cdot|w_{m}|$ becomes the new bias. We can thus reduce parameter count, since we can now store the kernels themselves along with a corresponding binary value indicating whether they are positive or negative kernels. Furthermore, each kernel can be neatly categorized into “contributing to good signal quality” (positive kernels) or “contributing to poor signal quality” (negative kernels). This process is performed after learning the kernels, with the goal of reducing the memory footprint and number of scalar operations for deployed models.
@@ 4.4 Correlated Kernel Pruning
Our kernel pruning method consists of the following steps:
@@
    Sort the kernel pairs by their cosine similarity (in order of descending similarity), and select the top $pa$ pairs to prune, where $pa$ is a hyperparameter.
@@
    For each selected pair to prune, determine which of the two kernels to remove and which to keep based on their effective contribution to the signal (described below).

### v4

# Sparse learned kernels for interpretable and efficient medical time series processing

@@ Abstract
Rapid, reliable, and accurate interpretation of medical time-series signals is crucial for high-stakes clinical decision-making. Deep learning methods offered unprecedented performance in medical signal processing but at a cost: they were compute-intensive and lacked interpretability. We propose Sparse Mixture of Learned Kernels (SMoLK), an interpretable architecture for medical time series processing. SMoLK learns a set of lightweight flexible kernels that form a single-layer sparse neural network, providing not only interpretability, but also efficiency, robustness, and generalization to unseen data distributions. We introduce a parameter reduction techniques to reduce the size of SMoLK’s networks while maintaining performance. We test SMoLK on two important tasks common to many consumer wearables: photoplethysmography (PPG) artifact detection and atrial fibrillation detection from single-lead electrocardiograms (ECGs). We find that SMoLK matches the performance of models orders of magnitude larger. It is particularly suited for real-time applications using low-power devices, and its interpretability benefits high-stakes situations.
@@ 1 Introduction
Medical signals such as electroencephalograms (EEGs), electrocardiograms (ECGs), and photoplethysmography signals (PPGs) have played a pivotal role in the field of medicine, serving as a crucial medium for diagnosing and monitoring patients’ conditions, facilitating timely interventions, and improving overall healthcare outcomes. These types of signals are usually manually reviewed by trained medical professionals, however, in today’s realm of endless medical data and the popularization of continuous passive monitoring, manual reviewing efforts have become inefficient or even a bottleneck in the patient care process. Traditional techniques to automate signal processing and categorization, while improving speed, often come with limitations in terms of performance. The rise of deep learning yielded a revolution in signal processing, offering models with unparalleled performance (e.g., [1, 2, 3, 4, 5]). However, these models often come with their own set of challenges, particularly in terms of computational demands and a lack of interpretability in their decision-making processes. Post-hoc saliency methods often fail to correctly identify feature importance for time-series signals from these black box models [6, 7].
Recognizing the need for high-performing interpretable models, we sought to design an architecture that combines the best of both worlds: the performance of deep learning with the computational efficiency and interpretability of traditional methods. We propose a method that has performance similar to – and often better than – state-of-the-art deep neural network (DNN) approaches but with several orders of magnitude fewer parameters. Our method, Sparse Mixture of Learned Kernels (SMoLK) learns a set of lightweight flexible kernels to construct a single-layer sparse neural network, providing a new efficient, robust, and interpretable medical signal processing architecture. To train these kernel functions, we introduce two parameter reduction techniques: weight absorption absorbs the kernel weighting factors into the kernels themselves and correlated kernel pruning reduces redundancies among the learned kernels.
@@
Our approach achieves compelling results in both the task of detecting artifacts in PPG signals, as well as the detection of atrial fibrillation from single-lead ECGs (an increasingly common task in modern-day wearables). For PPG artifact segmentation, SMoLK obtains the best of all worlds: (1) state-of-the-art (SotA) performance and (2) interpretability in a model with (3) several orders of magnitude fewer parameters than the current SotA. For the single-lead ECG task, SMoLK achieves metrics matching a deep ResNet with less than 1% the parameter count, while significantly outperforming deep networks in the low-data regime, even against equal-parameter deep networks controlling, for potential overfitting. For interpretability (2), SMoLK yields a tiny-yet-optimized set of convolution kernels that, when applied to the PPG signal, summed, and ceiled, directly produces a measure of PPG signal quality. This measure can then be thresholded to segment artifacts. The process is “nearly linear” given that the only non-linearity is a ceiling function setting negative values to zero. We can directly and precisely measure the contribution of any individual convolution to the predicted signal quality, and kernel contributions are directly proportional to the output signal. For classification tasks, we can directly invert the process to compute each part of the signal’s precise contribution to the final classification. Our approach is reminiscent of feature-detection-based approaches in that we learn a set of kernels that mimic handcrafted features. For our smallest model, we learn so few kernels (12 total) that we can inspect these features by eye and observe the waveforms learned by our approach. This contrasts with deep neural networks, where there are so many parameters one cannot feasibly inspect the inner workings of the model, and even if one took the time to meticulously inspect these features, it would still be unclear how each feature individually contributes to the output due to the built-in non-linearity.
In terms of efficiency and memory (3), our model reflects that for real-world applications, signal processing can be limited by computational resources or power consumption, as most wearable devices are small computing devices with limited battery life and low compute budget. Furthermore, memory is often a constraint as many wearable devices run multiple processes and monitor several metrics at once. SMoLK is simple to implement, and easily lend itself to memory and compute optimization. Our architecture can be quantized easily to 16-bit floating point operations without any appreciable decrease in performance. Our method is lightweight enough to allow real-time assessment of signal quality or classification of arrhythmias with a small enough impact that it could feasibly be implemented as a background process, using only kilobytes of memory.
@@ 2.2 Deep Learning Approaches
Most deep learning methods demand significant computational resources, rendering them impractical for use in wearable devices. An exception is Tiny-PPG [36], which emphasizes model pruning and the development of compact, portable models. Our experiments indicate that we achieve similar performance to Tiny-PPG but with much sparser models (about half the parameters). However, Tiny-PPG is not designed for interpretability. Trying to understand black box deep learning models is extremely challenging. Most work turns to methods that seek to explain black box model behavior, such as the aforementioned GradCAM and SHAP methods [42, 43], rather than build interpretability inherently into the model, which yields a faithful understanding of model behavior. Indeed, even modern explanation algorithms for medical signal processing with transformer architectures [45] fail to yield saliency maps that properly attribute features to their classification. In the aforementioned work, saliency maps for ST-elevation myocardial infarction correctly assign high saliency to some elevated ST segments, yet ignore equally prevalent ST segments seemingly at random. Furthermore, an area of equally intense saliency is assigned to a completely blank area of the ECG strip. Although it is not always obvious a priori what features are important, it is generally useful to know what features contributed to a model’s decision.
3 Model Architecture
@@ 3.1 Overview
To achieve this, we learn a set of $M$ convolution kernels of varying numbers to evaluate parameter scaling laws: $M=12$ (small), $72$ (medium), $384$ (large). The $M$ kernels are divided into sets of $M/3$ kernels of short (1.0 seconds), moderate (1.5 seconds), and long (3.0 seconds) sizes, with the goal of learning features of different lengths to improve classification quality. The kernels are convolved with the input PPG signal and the values are ceiled via $\max(0,x)$ to obtain a non-negative signal quality indicator. Additionally, a bias is added to each convolution channel. We compute a weighted sum of the convolutions to produce a single channel that represents a measure of signal quality, which is then transformed to range [0.0, 1.0] via the sigmoid function. We apply a 3rd-order Savitzky-Golay filter ($\sim$0.8-second filter window) to smooth the output, then apply thresholding at a value of 0.5 to produce a binary output segmentation. This pipeline is shown in Figure 1. The addition of a post-processing smoothing filter (Savitzky-Golay in our case) yields smoother segmentation maps and relies on the assumption that PPG artifacts tend to occur in groups, rather than as fragmented artifacts. However, this post-processing filter is not strictly necessary to maintain good DICE metrics and is an optional hyperparameter to be chosen for the use case (see Appendix A).
Figure 1: The PPG Processing Pipeline. The processing pipeline for the PPG signal quality segmentation task during training and during inference. First, a PPG signal is normalized to a unit normal by subtracting the mean and dividing by the standard deviation. Next, a set of convolutions is applied. Our smallest model is lightweight enough that all of the convolved signals are displayed in this figure. The convolved signals are ceiled, weighted, and summed, and a sigmoid is applied. Finally, the output is smoothed and thresholded to segment the signal. After training, similar kernels are ablated and kernel weights are absorbed to reduce parameter count.
@@
3.2 Classification
@@ 3.2.1 Overview
$\mathbf{f_{m}}(\mathbf{x})$ $=\text{max}(0,\mathbf{x}*\mathbf{k}_{m}+b_{m}).$
@@
$\phi_{m}(\mathbf{x})$ $=\text{Mean}(\mathbf{f_{m}}(\mathbf{x}))$
@@
$z_{j}(\mathbf{x})$ $=\sum_{m=0}^{M}w_{mj}\phi_{m}(\mathbf{x})+\sum_{\hat{f}\in\mathcal{F}}w_{\hat{f}j}\hat{f}(\mathbf{x})+b_{j}$
Where $\hat{f}(\mathbf{x})$ are the frequency bands of $\mathbf{x}$ in the power spectrum $\mathcal{F}$, and $w_{\hat{f}j}$ is the corresponding weight of that frequency band for class $j$. Finally, a non-linear activation function can be applied to this logit to generate a probability distribution. We use SoftMax in our case, which is defined as $\text{SoftMax}(z_{j})=\frac{e^{z_{j}}}{\sum_{{j^{\prime}}=1}^{n}e^{z_{j^{\prime}}}}$, where $z_{j}$ is the logit for class $j$. An illustrated overview of this pipeline can be seen in Figure 2a.
Figure 2: The ECG Processing Pipeline. a. The pipeline for processing a single lead ECG. First, a set of learned convolutions is applied to the ECG to produce several feature maps. Then, feature-wise means are computed, which are used as inputs to a linear model. Global frequency information, obtained from a power spectrum, is also used as inputs for the linear model to yield a logit. b. The reverse pipeline for interpreting the prediction in subfigure a from the learned kernel model. First, feature maps are generated by the learned kernel convolutions, in the same manner as a. However, instead of computing the mean value of this feature map, the feature maps are multiplied element-wise with their corresponding class weights. The power spectrum is not used in the interpretation. Lastly, a transposed convolution is applied to compute the importance of each part of the input signal. This method is almost an exact inverse of the forward procedure, allowing for a principled assignment of importance to the input signal that directly correlates to the output class logit. Notably, this method correctly labels the increased PR-interval as the most important feature in the classification of this ECG as $1^{st}$-degree heart block, indicating that our model has learned important features rather than spurious correlations.
@@ 3.2.2 Interpretability
Because our model is a linear combination of feature map mean activations, and each activation was generated from a convolution over the signal, we can directly compute the contribution of each part of the input signal to a given output class by reversing the convolution. We thus compute the contribution map as follows, in Algorithm 1. At a high level, a forward pass is performed as usual and the feature map values are computed. Then, the feature maps are multiplied element-wise with the feature weight for the class in question. Lastly, this product is convolved backward into a contribution map buffer of equal dimension to the input ECG. This allows us to compute exactly how much any given portion of the input signal contributed to the output classification. An illustration of this process can be seen in Figure 2b.
@@ 3.3 Weight Absorption
The quantity $\mathbf{k}_{m}\cdot|w_{m}|$ becomes the new kernel, and the quantity $b_{m}\cdot|w_{m}|$ becomes the new bias. We can thus reduce parameter count, since we can now store the kernels themselves along with a corresponding binary value indicating whether they are positive or negative kernels. Furthermore, each kernel can be neatly categorized into “contributing to good signal quality” (positive kernels) or “contributing to poor signal quality” (negative kernels). This process is performed after learning the kernels to reduce the memory footprint and number of scalar operations for deployed models.
@@ 3.4 Correlated Kernel Pruning
At a high level, our kernel pruning method consists of the following steps:
@@
    Sort the kernel pairs by their Euclidean distance (in order of increasing distance), and select the top $pa$ pairs to prune, where $pa$ is a hyperparameter.
@@
    For each selected pair to prune, determine which of the two kernels to remove and which to keep based on their effective contribution to the signal.
A rigorous description of the pruning process is discussed in the Methods section.
4 Methods
For all machine learning tasks, PyTorch 2.1.1 [54] and Scikit-learn 1.3.2 [55] were used. Data analysis and statistics were conducted with SciPy 1.9.3 [56] and NumPy [57] 1.22.4.
4.1 Datasets and Preprocessing
We train our model on the PPG-DaLiA [46] training split, and evaluate our approach on the same datasets used by the current state-of-the-art algorithms [41, 36], namely PPG-DaLiA [46] test split, WESAD [47], and TROIKA [48]. The PPG-DaLiA dataset is composed of recordings from 15 subjects performing real-world tasks. The data include electrocardiogram (ECG), accelerometer, and electrodermal recordings for various settings/activities, such as sitting still, walking up/down stairs, playing sports (soccer, cycling, walking), and driving. The WESAD data was recorded from wrist and chest PPGs from 15 subjects aged 21-55 (median 28) primarily focusing on varying the emotional state of the subject (neutral, stressed, amused). ECG, accelerometer, electrodermal, and wrist PPG signals were recorded. Lastly, the TROIKA data was recorded from 12 subjects running on a treadmill (ages 18-35). Accelerometer, ECG, and PPG signals were recorded. This dataset represents the most challenging dataset for generalization among the datasets we test in this work, since the subjects were under intense physical demands with poor signal quality, owing to the large amounts of movement.
4.2 PPG Segmentation
4.2.1 Training
4.2.2 Pruning